Find the product using suitable properties
854 × 102 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 854 × 102
= 854 × (100 + 2)
= 854 × 100 + 854 × 2 (using distributive property)
= 85400 + 1708
= 87108

Thanks